Sol Summary – February 7th

Crew 188 Sol 10 Summary Report 7FEB2018

Sol 10

Summary Title: Game On!
Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander Mission Status: Crew feeling good after a full day.

Sol Activity Summary:
Even with one EVA, the crew had a very busy and productive day. The EVA challenged the crew to a new slew of Martian Olympiad events counting down to the Winter Olympic Games back on Earth. An opening ceremony occurred and the crew hopes to share this artistic footage in high resolution 360 views. There are sure to be phots posted across social media, just look for #MDRS188. Meanwhile in the Hab, the NASA Glenn dust load experiment continued and air filtration dust samples were obtained from the EVA crew returning into the airlock. The afternoon was highlighted with staring at the Sun with a Hydrogen-alpha filter seeing sunspots, flares and even an airplane captured on camera passing in front of the Sun! The crew was busy digitally capturing video footage and audio for podcasts, school groups and for a mission documentary. After sunset the crew enjoyed watching the International Space Station pass-by briefly and an Iridium satellite. Who knows what space infrastructure Mars inhabitants will be able to watch from their homes, but they will need some form of planetary connectivity.

Spacesuit Up!
Ryan L. Kobrick, Ph.D.
MDRS Crew 188 Commander

Look Ahead Plan:
Two EVAs are panned that will fill most of the operational day. Afterwards the third and final human factors group chat will occur.

Anomalies in work:
Propane was detected in the afternoon. The crew is carefully monitoring and performing basic bubble testing on pipes. Currently no leaks are detected.

Weather:
Calm winds and clear skies with temperatures dropping rapidly at sunset.

Crew Physical Status:
One crewmember has cold symptoms, which may be due to fatigue but health status is nominal.

EVA:
Two are planned for tomorrow.

Reports to file:
1. Ops Report
2. Sol Summary
3. Journalist Report (combined Sol 9 and 10)
4. Green Hab Report
5. EVA #13 and #14 request
6. EVA #12 Reports
7. Daily Photos

Support Requested (same as previous report):
– Heater for science dome requested. A purchase order request was made.
– There are potentially two packages at Hollow Mountain that we would love to have picked up.
– Vegetables were requested due to the insufficient supply for six crewmembers. Non-vegetarians have cut their personal rations to help stretch the supplies for the two vegetarians. We have two celiac members who cannot eat the starches provided. These diets were all communicated before our arrival. In order to feed crewmembers over the next 4 days, 2-3 cans of dried vegetables were requested (or fresh) with the missing items from the original food supply. We were informed today that we will not be getting any form of resupply. Susan Holden Martin suggested we reiterate this request to mission support since it remains open. [end]

Operations Report – February 7th

Crew 188 Operational Report 07/02/2018

Sol 10

Zac Trolley

Non-Nominal Systems: electrical system, electrical cords

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 15.2h

Solar: SOC 78%

Diesel: 74%

Propane: 55 – Reading taken at 2:30 PM. There are no units on the gauge.

Ethanol Free Gasoline (5 Gallon containers for ATV): 3 Gallons

Water (trailer): 0 Gallons

Water (static): 120 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter 130875.1

Toilet tank emptied: No

ATVs Used: None

Oil Added: Yes, about 600ml total.

ATV Fuel Used: 0 Gallons

# Hours the ATVs were Used today: 0

Notes on ATVs: The Gasoline storage area needs to be moved. The sun is creating a large pressure differential on the jerry cans, making that area an explosion risk.

Deimos rover used: Yes

Hours: Hours: 108.1

Beginning charge: 100%

Ending charge: 100%

Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 18.0h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: Yes

Hours: 19.1h

Beginning charge: 100%

Ending charge: 99%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 11.6h

Beginning charge: 100%

Ending charge: 99%

Currently charging: No

HabCar: No

General notes and comments:

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Generator was turned off at 9:18

Generator was turned on at 18:20

I’ve found two extension cords that have been damaged and need to be replaced. If there are 3 prong plugs available, I can repair the cords. Please advise.

The propane smell was back today. Around 5:00pm there was two people in the Hab, and the smell was detected in the same place as before. The downstairs heater was off. It has been my understanding that the downstairs heater turns off automatically when it reaches a certain temperature. The source of the gas could be a portion of propane that is not combusted when the heater automatically turns off.

If this is NOT the way the heater functions, please let me know.

We used the detector and there was a positive detection of gas. Unfortunately, we did not think to do a control check with the instrument when no gas was smelt to ensure we are not getting false positives. I did another bubble test on all the visible gas connections, and nothing was found.

Summary of internet: Tests were run on the internet connectivity and passed on to Bernard.

Summary of suits and radios: Nominal

Summary of Hab operations: We are running low of vegetables, and there is a concern that our vegetarians will not have enough food for the remainder of the mission.

Summary of EVA operations: Nominal

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: We could use another bag of toilet deodorizer, as we only have one puck left.

EVA Report – February 6th

Crew 188 EVA #10 Summary Report 06FEB2018
Author’s name: Julia DeMarines, Crew Astronomer, Green Hab Operator

Purpose of EVA: Filming of artistic Performing Astronautics and Bending Horizons projects using the Insta360 Pro Camera.

Location of EVA: RC South Site Region
UDM27 Coordinates: 12 S 518293 E 4249791 N (waypoint taken on EVA009)

Number of EVA Crew: 3
Participants: Zac Trolley, Sarah Jane Pell, Julia DeMarines
EVA Commander: Sarah Jane Pell

Road(s) and routes per MDRS map:
1. Take East on Entrance Road (Driving)
2. Turn Right onto Cow Dung Road (Driving)
3. Park on Cow Dung Road behind filming area ridge
4. Walk to filming location in field

Mode of Travel: Rover and Pedestrian
Vehicles you will be using (if applicable): Curiosity and Spirit

Summary:

EVA #10 arrived at our destination to the filming location on a very windy morning. We scouted the area out making sure this is where we wanted to film and if we could find a more wind-protected area to shield the Insta360º Pro camera. After some scouting we decided that the original location was ultimately the best location for Sarah Jane’s performance art. We discussed the choreography, set up the camera and then began recording. The first take was a success!

After completing Sarah Jane’s Performing Astronautics piece recorded using the Insta360º Pro, lovingly called Magpie, EVA 10 performed a science experiment in honor of the Falcon Heavy launch today. Julia acquired two small pop-top containers (that were once used for electrolyte tablets) and several alka-seltzer tablets. When small chunks of the alka-seltzer tablets are mixed with water in a closed container, gas builds up as does pressure. It took a few tries to get the ratio of water, air, and tablets correct, but once we did, we were able to launch our “Falcon Light” mini rockets! Our landing was not as graceful as the Falcon Heavy’s boosters returning in sync but we do believe our successful launch was 100% responsible for the success of the launch today as pop-rockets are known for their good luck and scientific opulence.

We will send photos and videos tomorrow.

EVA Report – February 6th

Crew 188 EVA #11 Summary Report 6FEB2018

Author’s name: Tatsunari Tomiyama AHFP, Health & Safety Officer Purpose of EVA: Reconnaissance of “Moon” region near Brache Hwy 1572 for future 360 filming and micrometeorite samples in vicinity nears roads travelled.
Location of EVA: Copernicus Highway 1574 waypoints and Skyline Ridge as time permits.
UDM27 Coordinates: 12 S 515404E, 4253907N

Number of EVA Crew: 3
Participants: Ryan Kobrick Tatsunari Tomiyama, Renee Garifi
EVA Commander: Ryan Kobrick

Rovers used: ATV 1, 2, 3
Duration: 2hours 23minutes
EVA Departure Time: 14:03
EVA Return Time: 16:26

Summary:
The objective of this EVA was to create EVA way-points for a future mission. We started to prepare EVA #11 before the EVA #10 group arrived back to the base to maximize EVA activity time because we expected that the destination point could be far away than our planning expectation. We used 3 ATVs for this reconnaissance mission. Brahe Highway was relatively easy to navigate using GPS but has many large bumps and pits in the road and required us to go slowly during this part of the route. However, once people have experience, this issue could be negligible. Overall, the objective was successfully completed in this EVA activity although our Garmin GPS unit battery was running low due to use on both EVA #10 and EVA #11.

Following our micrometeorite sample collections on a flat hilltop, we took Brahe Highway to the right on the Copernicus Highway and headed to east. We passed “Yellow Moon” indicated on the MDRS provided map. At that point, we returned back the Copernicus highway toward west. We could drive nearby Toothy Ridge on the Copernicus Highway and we decided to return to the base because of the time restriction.

On the way back to the base, we passed a grey unmarked electric rover parked about 4 minutes north of the Hab on Cow Dung Road. The vehicle was the same color and the same type which we use for MDRS. There was no person around the vehicle and no name was labeled on the vehicle. We took a photo in case mission support would like to see the vehicle. It may be the personal rover of Dr. Rupert but we would like to confirm.

Sol Summary – February 6th

Crew 188 Sol 9 Summary Report 6FEB2018

Sol 9

Summary Title: Sacrifice
Author’s name: Ryan L. Kobrick, Ph.D., MDRS Crew 188 Commander Mission Status: Crew is well.

Sol Activity Summary:

I think it goes without saying that our crew made many sacrifices to be a part of this two-week experience (finances, family, work, and much more). One that may not be obvious, and would have been impossible to time, was missing the Falcon Heavy historic launch. The unquestionable highlight for today was SpaceX’s success launch and landings, and what that may mean for future human spaceflight exploration. Today was packed with exploration. Two EVAs including a beautifully choreographed performance optimizing a 360 camera, and a driving excursion to uncharted territory. The crew was able to watch the launch and twin landings in a video later in the evening. Wow. Even watching the launch delayed had a similar impact to being there with family and friends in Florida. It was rocket-eye-candy-overload. Who knows when the next FH launch will be, or where any of us may be located, but we do know that these are the ingredients needed to make a mission to Mars feasible.

Spacesuit Up!
Ryan L. Kobrick, Ph.D.
MDRS Crew 188 Commander

Look Ahead Plan:
Tomorrow solar observations will continue in the morning hoping for calmer weather. An afternoon EVA will celebrate our international diversity and the upcoming Olympic Games with our own Martian Olympiad.

Anomalies in work:
None.

Weather:
The temperatures are cooling off and the winds remain high.

Crew Physical Status:
The crew has been burning on full throttle from 8am until 11pm and pushing hard to the finish line to be fully successful in their scientific and artistic investigations at MDRS.

EVA:
One is planned for tomorrow.

Reports to file:
1. Ops Report
2. Sol Summary
3. Journalist Report
4. Green Hab Report
5. EVA #12 request
6. EVA #10 and #11 Reports
7. Daily Photos

Support Requested:
– Heater for science dome requested. A purchase order request was made.
– There are potentially two packages at Hollow Mountain that we would love to have picked up.
– Vegetables were requested due to the insufficient supply for six crewmembers. Non-vegetarians have cut their personal rations to help stretch the supplies for the two vegetarians. We have two celiac members who cannot eat the starches provided. These diets were all communicated before our arrival. In order to feed crewmembers over the next 4 days, 2-3 cans of dried vegetables were requested (or fresh) with the missing items from the original food supply. We were informed today that we will not be getting any form of resupply. Susan Holden Martin suggested we reiterate this request to mission support since it remains open.
– Our reports have not been posted to the MDRS website since Saturday Feb 3rd. Please check the status of these documents for dissemination. We did see some photos on Facebook.

Greenhab Report – Frebruary 6th

Crew 188 Sol 9 Green Hab Report 06FEB2018

Zac Trolley (Engineer), Guest writing for Julia De Marines (Crew Astronomer/GreenHab Officer)

06 February 2018

Environmental control: Heating

Ambient with door opening: Shade cloth on

Working hours in Green Hab: 11:00 – 11:30, 18:00 – 19:00

Outside temp at working hour: 15ºC

Inside temp at working hour: 26ºC

Inside temp H: 33ºC

Inside temp L: 14ºC

Inside humidity: 23 %

Inside humidity H: 81%

Inside humidity L: 19%

Hours of supplemental light: 05:00 – 9:59

Changes to the crops: We picked many beans, herbs, and green leaves to dine upon.

Daily water usage for crops: 48 litres

Time(s) of watering plants: 11:00 – 11:30 (16 litres), 18:00 – 19:00 (32 litres)

Changes to research plants: The VEGGIE OASYS is progressing.

Narrative: Tat, our intrepid health and safety officer bravely foraged among the forest of plants being nurtured by our crew to bring home a bounty of green beans. Sadly, Sarah Jane has been unable to bridge the plant-human communication barrier.

​Future needs and questions: We are awaiting the arrival of the tomato cages so we can replant the tall plants that require support to grow to their full potential. Please advise when they will arrive. ​

Operations Report – February 6th

Crew 188 Operational Report 06/02/2018

Sol 9

Zac Trolley

Non-Nominal Systems: electrical system, electrical cords

Notes on non-nominal systems: See individual systems at the end of the report.

Generator (hours run): 15.2h

Solar: SOC 74%

Diesel: 75%

Propane: 61 psi

Ethanol Free Gasoline (5 Gallon containers for ATV): 3 Gallons

Water (trailer): 0 Gallons

Water (static): 170 Gallons

Trailer to Static Pump used: No

Water (loft) – Static to Loft Pump used: Yes

Meter 130825.4

Toilet tank emptied: Yes

ATVs Used: The three blue ATVs were used today

Oil Added: None

ATV Fuel Used: 2 Gallons

# Hours the ATVs were Used today: 2

Notes on ATVs: It is very difficult to see the fuel level in the ATVs with the suit helmet on. Also, the gas can that had a little fuel in the bottom imploded due to temperature change. There was no leak, but the structural integrity of the can has been compromised and it should be destroyed. I suggest a new location be found to keep the cans out of the sun.

Deimos rover used: No

Hours: Hours: 108.1

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: No

Hours: 18.0h

Beginning charge: N/A

Ending charge: N/A

Currently charging: No

Opportunity rover used: Yes

Hours: 19.0h

Beginning charge: 100%

Ending charge: 97%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 11.4h

Beginning charge: 100%

Ending charge: 99%

Currently charging: No

HabCar: Yes, to get a load of water.

General notes and comments:

During an EVA, a plastic handle from the passenger side roll cage on Demos came off. We have the handle, but no hardware to re-attach it.

Generator was turned off at 9:00

Generator was turned on at 18:04

I’ve found two extension cords that have been damaged and need to be replaced. If there are 3 prong plugs available, I can repair the cords. Please advise.

Summary of internet: Tests were run on the internet connectivity and passed on to Bernard.

Summary of suits and radios: Nominal

Summary of Hab operations: We are running low of vegetables, and there is a concern that our vegetarians will not have enough food for the remainder of the mission.

Summary of EVA operations: Nominal

Summary of GreenHab operations: The temperature gauge display has some non-functional LCD areas, a new one has been ordered.

Summary of Science Dome operations: The science dome is awaiting a portable heater as per the MDRS Director.

Summary of RAM operations: Not Operational.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: We could use another bag of toilet deodorizer, as we only have one puck left.

The “tunnel” had a few sharp edges of the metal at the corners, and a jacket was torn. I’ve taped over the edges so it doesn’t happen again.

Greenhab Report – February 5th

Crew 188 Sol 7 Green Hab Report 05FEB2018

Zac Trolley (Engineer), Guest writing for Julia De Marines (Crew Astronomer/GreenHab Officer)

05 February 2018

Environmental control: Heating

Ambient with door opening: Shade cloth on

Working hours in Green Hab: 11:00 – 11:30, 18:00 – 19:00

Outside temp at working hour: 17ºC

Inside temp at working hour: 25ºC

Inside temp H: 35ºC

Inside temp L: 14ºC

Inside humidity: 27 %

Inside humidity H: 79%

Inside humidity L: 18%

Hours of supplemental light: 05:00 – 9:59

Changes to the crops: The plants are following a predictable pattern of requiring sun and water, and producing green leaves. As the plants increase in size, we have been harvesting various parts of them to eat in our kitchen. On this 8th Sol, we dined on lettuce leaves and tasty herbs.

Daily water usage for crops: 48 litres

Time(s) of watering plants: 11:00 – 11:30 (16 litres), 18:00 – 19:00 (32 litres)

Changes to research plants: During an intensive research investigation it was discovered that one of the pillows in the VEGGIE OASYS experiment isn’t as wet as is desired. Further investigation is required in order to find the route cause of this discrepancy. The Executive Officer is working with the NASA PI in order to properly moisten the pillow.

Narrative: In order to better understand the needs and desires of the plants under our care, we have been closely monitoring their progress. We have been exploring alternative methods to increase plant yield, and have decided to ask the plants what they require to be at their maximum efficiency. To facilitate this, Zac Trolley and Sarah Jane Pell have undertaken an experimental procedure to breach the communication gap with the plants and create a dialogue with them directly. So far, our results are not promising.

Future needs and questions: We are awaiting the arrival of the tomato cages so we can replant the tall plants that require support to grow to their full potential. Please advise when they will arrive.

Sol Summary – February 5th

Crew 188 Sol 8 Summary Report 05FEB2018

Sol 8

Summary Title: Team Unity
Author’s name: Renee Garifi, MDRS Crew 188 Executive Officer Mission Status: Nominal

Sol Activity Summary:
This morning our team awoke to a beautifully and thoughtfully arranged breakfast table with a special greeting written by two crewmembers who wanted to help our team start week 2 of our mission on a high note. Each mission brings with it a unique set of challenges and rewards and I am humbled and proud to be part of such a courageous and strong team.

Every morning at 08:30 we hold our Daily Planning Conference and one crewmember hosts a briefing of the upcoming EVA or team activity that will require the most coordination by all crew. We spent the mid-morning preparing for a 4-person EVA with the objective of scouting filming locations for a performing astronautics project by our Artist in Residence utilizing the Insta360 Pro camera. Two of our crew remained behind at the Hab to serve as HABCOM on the radio with the EVA team and keep up with science projects and tasks around the Hab.

Following the EVA we shared lunch and set about working on our individual projects. We had a productive afternoon followed by a period of relaxation. Our crew is currently in good spirits and communication, stress management and team work is at an all-time high. In the evenings when our crew is not too exhausted, we have started hosting a lecture series here on Mars. After our comms window closes and our crew finishes dinner and dishes, the next scheduled crewmember to present will spend about 30-45 minutes sharing a topic that they are either passionate about or teaching a useful skill. Our crew is comprised of six individuals with strong personal and professional backgrounds that make for informative evening sessions that bond and inspire us.

Ad Astra!
Renee Garifi
MDRS Crew 188 Executive Officer

Look Ahead Plan:
We will perform EVA #10 in the morning tomorrow and EVA #11 in the afternoon.

Anomalies in work:
None.

Weather:
Windy, but not like the high winds of yesterday. Skies were cloudy.

Crew Physical Status:
The crew is tired but in good spirits.

EVA:
Two EVAs planned for tomorrow.

Reports to file:
1. EVAs #10 and #11 request
2. EVA #9 Report
3. Ops Report
4. Sol Summary
5. Journalist Report
6. Green Hab Report
7. Daily Photos

Support Requested (acknowledged yesterday but listed as reminders): – Heater for science dome requested.
– There are potentially two packages at Hollow Mountain that we would love to have picked up.
– Water tank refills will be needed in the next few days.